You will need to calculate your Maximal and Training Heart Rate Ranges, be sure to show...

  1. You will need to calculate your Maximal and Training Heart Rate Ranges, be sure to show your work.
    1. Maximal Heart Rate- 220- age= Max beats per minute
    2. Lower end of intensity range 50% of max (0.50x Max= low intensity)
    3. Moderate intensity 70% of max (0.70 x Max= moderate intensity)
    4. Vigorous intensity 85% of max (0.85 x Max = vigorous intensity

(ENTER AN IDEAL HEART RATE TO WORK WITH IF REQUIRED)
Answer #1

Age: - 20 years  

Normal Heart rate: 78 beats / min

1. Maximum bets / min = 220- age

                                    = 220-20

                                      = 200 Beats / min

2. Low end of intensity range 50 % of maximum beats / min

                                    = 0.50 * Max beats / min

                                    = 0.50*200

                                    = 100 beats / Min

3. Moderate Intensity = 0.70 * Max beats / Min

                                    = 0.70*200

                                    = 140 beats / Min

4. Vigorous intensity = 0.85 * Max beats / Min

                                    = 0.85 * 200

                                    = 170 beats /min